OpenStudy (anonymous):

Solve for k. 8k + 2m = 3m + k k = 70m k = 7m k = 7/m k = m/7

OpenStudy (liv1234):

Do you know how to solve these?

OpenStudy (anonymous):

no because havent done algebra since 9thgrade

OpenStudy (gabebae):

Combine like terms

OpenStudy (anonymous):

sorry dont know how to

OpenStudy (gabebae):

\[8k + 2m = 3m + k\] subtract -2m from both sides \[8k = m + k\] Subtract k from both sides \[7k = m\] Divide both sides by 7 \[k = \frac{ m }{ 7 }\]

OpenStudy (anonymous):

oh ok thnx so much

OpenStudy (gabebae):

you're welcome